OpenStudy (anonymous):

what are the zeros of the polynomial function f(x) = x^3 – x^2 - 12x

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@ybarrap @wild_rebel_gurl21

OpenStudy (anonymous):

@coolsday

OpenStudy (anonymous):

lol....thanks I guess. What is it that you need? :)

OpenStudy (anonymous):

Well I would like to know how to find my zeros for my problem up above!

OpenStudy (anonymous):

factor it?

OpenStudy (anonymous):

Five

OpenStudy (anonymous):

what do u mean 5?

OpenStudy (anonymous):

my options are like 0,4,3 or 0,-4,3

OpenStudy (anonymous):

etc like that

OpenStudy (anonymous):

The formula is written wrong.

OpenStudy (anonymous):

fx is not obsolute equation

OpenStudy (anonymous):

well thats how they give it to me

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

x^3 – x^2 - 12x=x(x^2-x-12)=x(x-4)(x+3) Set each factor equal to 0 x = 0 x-4=0 x+3=0 Solve each equation x=0 x=4 x=-3

OpenStudy (anonymous):

Given polynomial f(x) if (x+7) is a factor what else must be true

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@Nikkeyiboo

OpenStudy (mertsj):

If x+7 is a factor then x=-7 must be a root.

OpenStudy (anonymous):

f(0) = -7?

OpenStudy (anonymous):

?????

OpenStudy (mertsj):

@crystelle That would be f(-7)=0